# FedOne: Query-Efficient Federated Learning for Black-box Discrete Prompt Learning

### v1

@@ 2.1 Federated Black-box Prompt Learning Framework
[…]
@@ 2.2 Black-box Discrete Prompt Learning on the Client
^$\quad\ \hat{\nabla}_{\boldsymbol{\alpha}_{i}^{k}}f^{k}(\boldsymbol{\alpha}^{k},\mathcal{B}^{k})$
@@ 2.3 Algorithm
18:     Compute ^$\hat{\nabla}_{\boldsymbol{\alpha}_{i}^{k}}f^{k}(\boldsymbol{\alpha}^{k},\mathcal{B}^{k}_{e})$ using Eq. 2.2.
@@ 3.1 Assumptions
Bounded Clients’ Heterogeneity: For client $k=1,...,K$ with sampling probability vector $\boldsymbol{q}=\left\{q^{\left[k\right]}\right\}_{k=1}^{K}$, and $\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k},\Psi^{k})$ is local gradient of $\boldsymbol{\alpha}_{i}$ w.r.t. all input sentence in client $k$. We assume that $\lambda$ is the upper bound on the weighted gradient diversity across local objectives, i.e.
@@ 3.2 The Convergence of Federated BDPL
and establish that the minimum of $Q_{\epsilon}(K_{*})$ is achieved when $10<K_{*}<$ 1. Consequently, $Q_{\epsilon}(K_{*})$ is a monotonically increasing function within the feasible region of $K_{*}$, for $K_{*}=1,2,\dots,K$. Thus, to minimize query overhead and achieve optimal efficiency, the optimal choice is $K_{*}=1$.
@@ 4.1 Experiment Setup
The model architecture employed is RoBERTa-large (Liu et al., 2019). The trainable prompts were placed at different positions in the model depending on the algorithm of the baselines. For the training procedure, we conducted a hyperparameter tuning phase using a grid search approach to explore learning rates of $[3e-4,1e-4,3e-5,1e-5]$. The batch size was set at 32, and the optimization algorithm employed was AdamW (Loshchilov & Hutter, 2017). Further details about the dataset and evaluation metrics of them are available in Appendix B.1.
@@ 4.3 Result
Baseline Comp. Time for Training (s) FL Server Comm. Cost (MB) FL Server # Queries LLM Server # Query Per Client Trainable Parameter Size (MB) Per Client Loaded GPU Memory (MB)
[…]
@@
No Prompt ${14.05}_{0.00}$ ${68.04}_{0.00}$ ${91.35}_{0.00}$ ${79.62}_{0.00}$ ${36.01}_{0.00}$ ${56.22}_{0.00}$ ${72.20}_{0.00}$ 59.6459.64{59.64}59.64
Prompt w/o. Training ${9.17}_{1.31}$ ${62.94}_{4.94}$ ${79.47}_{2.71}$ ${79.88}_{1.94}$ ${25.53}_{1.10}$ ${56.69}_{4.76}$ ${72.16}_{2.51}$ 54.4154.41{54.41}54.41
FedOne-BDPL $13.67_{0.17}$ $68.25_{3.05}$ $87.77_{3.82}$ $81.69_{0.70}$ $32.16_{6.51}$ $70.58_{1.35}$ $77.62_{1.62}$ 61.67
@@ A.1 Notations and Omitted Mathematical Steps
$t=0,...,T-1$ Number of global iterations
$e=1,...,E$ Number of local iterations
$s=0,...,S$ Number of client-server interactions
[…]
@@
^$\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\cdot,\cdot)$ The local stochastic mini-batch variance-reduced policy gradient of the client $k$ (2.2)
@@
^$\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{\ast}(\cdot,\cdot)$ The average stochastic mini-batch variance-reduced policy gradient of the client group $U_{t}$ (92)
@@
For $i=1,...,n$, we define the stochastic gradient, stochastic mini-batch gradient and full gradient with respect to $\boldsymbol{\alpha}^{k}_{i}$ as following:
@@ A.3 Lemmas
Unbiasedness and bounded variance of variance-reduced policy gradient: At the $k$-th client, $\forall\ \psi_{m}^{k}\in\Psi^{k}$, $r=1,\cdots,I$ denotes the $r$-th sampling of $\Phi^{k}$ from $\boldsymbol{p}^{k}$ w.r.t. $\left\{\Phi^{k,r}\right\}_{r=1}^{I}\sim\boldsymbol{p}^{k}$, $\boldsymbol{p}^{k}=\text{GS}(\boldsymbol{\alpha}^{k})$, $\alpha_{i,j}\geq\nu>0$ for $i=1,...,n$ and $j=1,...,N$, $\tau>0$ is the temperature parameter, and $\sigma_{\alpha}^{2}=\frac{8G^{2}N}{\tau^{2}\nu^{2}}$, then the variance-reduced policy gradient is unbiased and bounded by :
@@
=(1)$\overset{(1)}{=}\frac{1}{I}\sum_{r=1}^{I}\mathbb{E}_{\Phi^{k,r}}\left[\mathcal{L}(\Phi^{k,r},\psi_{m}^{k})\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}\log P(\phi_{i}^{k,r})\right]$
@@
=$\overset{(2)}{=}\nabla_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k},\psi_{m}^{k})-\frac{1}{I}\sum_{r=1}^{I}\left(\frac{1}{I-1}\sum_{\begin{array}[]{c}w=1\\ w\neq r\end{array}}^{I}\mathbb{E}_{\Phi^{w}}\left[{\mathcal{L}}(\Phi^{k,w},\psi_{m}^{k})\right]\right)\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}\left(\sum_{\phi_{i}^{k,r}\sim\boldsymbol{p}^{k}_{i}}P(\phi_{i}^{k,r})\right)$ (46)
=$\overset{(3)}{=}\nabla_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k},\psi_{m}^{k})-\frac{1}{I}\sum_{r=1}^{I}\left(\frac{1}{I-1}\sum_{\begin{array}[]{c}w=1\\ w\neq r\end{array}}^{I}\mathbb{E}_{\Phi^{w}}\left[{\mathcal{L}}(\Phi^{k,w},\psi_{m}^{k})\right]\right)\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}\left(1\right)$ (49)
@@
=(1)$\overset{(1)}{=}\frac{1}{I^{2}}\sum_{r=1}^{I}\mathbb{E}_{\Phi^{k,r}}\left[\left\|\frac{1}{I-1}\sum_{\begin{array}[]{c}w=1\\ w\neq r\end{array}}^{I}\left(\mathcal{L}(\Phi^{k,r},\psi_{m}^{k})-{\mathcal{L}}(\Phi^{k,w},\psi_{m}^{k})\right)\nabla_{\boldsymbol{\alpha}^{k}_{i}}\log P(\phi_{i}^{k,r})-\nabla_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k},\psi_{m}^{k})\right\|^{2}\right]$ (55)
@@
L-smooth for $\boldsymbol{\alpha}^{k}$: At the $k$-th client, the $\Phi^{k}$ is sampled from probability matrix $\boldsymbol{p}^{k}$, and $\boldsymbol{p}^{k}=\text{GS}(\boldsymbol{\alpha}^{k})$, $\alpha_{i,j}\geq\nu>0$ for $i=1,...,n$ and $j=1,...,N$, $\tau>0$ is the temperature parameter, $\mathbb{E}_{\Phi^{k}\sim\text{GS}(\boldsymbol{\alpha}^{k})}\left[{\mathcal{L}}(\Phi^{k},\Psi^{k})\right]$ is L-smooth for $\boldsymbol{\alpha}^{k}$ and $L=\frac{nGN(\tau+1)}{\tau^{2}\nu^{2}}$, and then for $t$-th iteration, the following inequality is satisfied:
@@
$\nabla_{\boldsymbol{\alpha}^{k}}\mathbb{E}_{\Phi^{k}_{t}\sim\text{GS}(\boldsymbol{\alpha}^{k}_{t})}\left[{\mathcal{L}}(\Phi^{k}_{t},\Psi^{k})\right],\boldsymbol{\alpha}_{t+1}^{k}-\boldsymbol{\alpha}_{t}^{k}\right>+\frac{L}{2}\left\|\boldsymbol{\alpha}_{t+1}^{k}-\boldsymbol{\alpha}_{t}^{k}\right\|^{2}$
@@
$\frac{\partial^{2}\left[P(\phi_{i^{\prime}}^{k})P(\phi_{i^{\prime\prime}}^{k})\right]}{\partial\alpha_{i^{\prime},j^{\prime}}\partial\alpha_{{i}^{\prime\prime},{j}^{\prime\prime}}}=\begin{cases}\frac{1-p_{i^{\prime},j^{\prime}_{i}}^{k}}{\tau\alpha_{i^{\prime},j^{\prime}_{i}}^{k}}\cdot\frac{1-p_{i^{\prime\prime},j^{\prime\prime}_{i}}^{k}}{\tau\alpha_{i^{\prime\prime},j^{\prime\prime}_{i}}^{k}},\text{ if }j^{\prime}=j^{\prime}_{i}\text{ and }j^{\prime\prime}=j^{\prime\prime}_{i}\\ \frac{1-p_{i^{\prime},j^{\prime}_{i}}^{k}}{\tau\alpha_{i^{\prime},j^{\prime}_{i}}^{k}}\cdot\left(-\frac{p_{i^{\prime\prime},j^{\prime\prime}}^{k}}{\tau\alpha_{i^{\prime\prime},j^{\prime\prime}}^{k}}\right),\text{ if }j^{\prime}=j^{\prime}_{i}\text{ and }j^{\prime\prime}\neq j^{\prime\prime}_{i}\\ -\frac{p_{i^{\prime},j^{\prime}}^{k}}{\tau\alpha_{i^{\prime},j^{\prime}}^{k}}\cdot\frac{1-p_{i^{\prime\prime},j^{\prime\prime}_{i}}^{k}}{\tau\alpha_{i^{\prime\prime},j^{\prime\prime}_{i}}^{k}},\text{ if }j^{\prime}\neq j^{\prime}_{i}\text{ and }j^{\prime\prime}=j^{\prime\prime}_{i}\\ -\frac{p_{i^{\prime},j^{\prime}}^{k}}{\tau\alpha_{i^{\prime},j^{\prime}}^{k}}\cdot\left(-\frac{p_{i^{\prime\prime},j^{\prime\prime}}^{k}}{\tau\alpha_{i^{\prime\prime},j^{\prime\prime}}^{k}}\right),\text{ if }j^{\prime}\neq j^{\prime}_{i}\text{ and }j^{\prime\prime}\neq j^{\prime\prime}_{i}\end{cases}$ (75)
@@
Convergence of BDPL: Suppose Assumption 3.1 and 3.2 hold, for $t=0,...,T-1$, $\alpha_{i,j}\geq\nu>0$ for $i=1,...,n$ and $j=1,...,N$, $\tau>0$ is the temperature parameter, $\mathbb{E}_{\Phi^{k}\sim\text{GS}(\boldsymbol{\alpha}^{k})}\left[{\mathcal{L}}(\Phi^{k},\Psi^{k})\right]$ is L-smooth for $\boldsymbol{\alpha}^{k}$ and $L=\frac{nGN(\tau+1)}{\tau^{2}\nu^{2}}$, $\sigma_{\psi}^{2}$ is the variance of the stochastic gradient, $\sigma_{\alpha}^{2}=\frac{8G^{2}N}{\tau^{2}\nu^{2}}$ is the variance of the variance-reduced policy gradient. Let $\eta\leq\frac{1}{L}$, then the BDPL’s full gradient $\nabla F^{k}(\boldsymbol{\alpha}_{t}^{k},\Psi^{k})$ satisfies the following inequality:
@@
$\nabla_{\boldsymbol{\alpha}^{k}}\mathbb{E}_{\Phi_{t}^{k}\sim\text{GS}(\boldsymbol{\alpha}_{t}^{k})}\left[{\mathcal{L}}(\Phi_{t}^{k},\Psi^{k})\right],\boldsymbol{\alpha}_{t+1}^{k}-\boldsymbol{\alpha}_{t}^{k}\right>+\frac{L}{2}\left\|\boldsymbol{\alpha}_{t+1}^{k}-\boldsymbol{\alpha}_{t}^{k}\right\|^{2}$
@@
=$\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{t},\Psi^{k}),-\eta\cdot\mathbb{E}_{\{\Phi^{r}\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}}\left[\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{t},\mathcal{B}^{k}_{t})\right]\right>+\frac{L\eta^{2}}{2}\mathbb{E}_{\{\Phi^{r}\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}}\left[\left\|\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{t},\mathcal{B}^{k}_{t})\right\|^{2}\right]$
=$\overset{(1)}{=}\left<\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{t},\Psi^{k}),-\eta\cdot\mathbb{E}_{\mathcal{B}_{t}}\mathbb{E}_{\{\Phi^{r}\}_{r=1}^{I}}\left[\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{t},\mathcal{B}^{k}_{t})\right]\right>$
@@
=$\overset{(2)}{=}-\eta\cdot\left<\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{t},\Psi^{k}),\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{t},\Psi^{k})\right>+\frac{L\eta^{2}}{2}\left\|\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{t},\Psi^{k})\right\|^{2}$
@@
[…]

### v2

@@ 2.1 Federated Black-box Prompt Learning Framework
$\quad\text{where:~ }{\mathcal{L}}^{k}(\Phi;\Psi^{k})=\frac{1}{M^{k}}\sum_{m=1}^{M^{k}}\ell\left(\Phi;\psi^{k}_{m},y_{m}^{k}\right)$ (1)
@@ 2.2 Black-box Discrete Prompt Learning on the Client
$\quad\ \hat{\nabla}_{\boldsymbol{\alpha}_{i}^{k}}f^{k}(\boldsymbol{\alpha}^{k},\mathcal{B}^{k})$
@@ 2.3 Algorithm
18:  Compute $\hat{\nabla}_{\boldsymbol{\alpha}_{i}^{k}}f^{k}(\boldsymbol{\alpha}^{k},\mathcal{B}^{k}_{e})$ using Eq. 2.2.
@@ 3.1 Assumptions
Bounded Clients’ Heterogeneity: For client k=1,,Kk=1,...,K with sampling probability vector $\boldsymbol{q}=\left\{q^{\left[k\right]}\right\}_{k=1}^{K}$, and $\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k},\Psi^{k})$ is local gradient of $\boldsymbol{\alpha}_{i}$ w.r.t. all input sentence in client $k$. We assume that $\lambda$ is the upper bound on the weighted gradient diversity across local objectives, i.e.
@@ 3.2 The Convergence of Federated BDPL
and establish that the minimum of $Q_{\epsilon}(K_{*})$ is achieved when $0<K_{*}<1$. Consequently, $Q_{\epsilon}(K_{*})$ is a monotonically increasing function within the feasible region of $K_{*}$, for $K_{*}=1,2,\dots,K$. Thus, to minimize query overhead and achieve optimal efficiency, the optimal choice is $K_{*}=1$.
@@ 4.1 Experiment Setup
The model architecture employed is RoBERTa-large (Liu et al., 2019). The trainable prompts were placed at different positions in the model depending on the algorithm of the baselines. For the training procedure, we conducted a hyperparameter tuning phase using a grid search approach to explore learning rates of [3e4,1e4,3e5,1e5][3e-4,1e-4,3e-5,1e-5]. The batch size was set at 32, and the optimization algorithm employed was AdamW (Loshchilov & Hutter, 2017). Further details about the dataset and evaluation metrics of them are available in Appendix B.1.
@@ 4.3 Result
Baseline
Comp. Time
for Training (s)
FL Server
Comm. Cost
FL Server
# Queries
LLM Server
# Query
Per Client Trainable
Parameter Size
Per Client Loaded
GPU Memory (MB)
Fed-Prompt-Tuning 224.26{224.26} 156.25{156.25} 1000{1000} - $7.8\times 10^{-2}$} $3564\text{ (Model, Grad., Prompt)}$}
Fed-P-Tuning v2 276.23{276.23} 11417.20{11417.20} 1000{1000} - 5.7{5.7} $3656\text{ (Model, Grad., Prompt)}$}
Fed-BBT 24501.83{24501.83} 3.81{3.81} 1000{1000} 200000{200000} $1.9\times 10^{-3}$} $1.90\times 10^{-3}\text{ (Prompt Only)}$}
Fed-BDPL 2407.05{2407.05} 30.52{30.52} 1000{1000} 20000{20000} $1.5\times 10^{-2}$} $1.52\times 10^{-2}\text{ (Prompt Only)}$}
Fed-GS-BDPL 2400.60{2400.60} 30.52{30.52} 1000{1000} 20000{20000} $1.5\times 10^{-2}$} $1.52\times 10^{-2}\text{ (Prompt Only)}$}
FedOne-Prompt-Tuning 16.99{16.99} 15.63{15.63} 100{100} - $7.8\times 10^{-2}$} $3564\text{ (Model, Grad., Prompt)}$}
FedOne-P-Tuning v2 19.95{19.95} 1141.72{1141.72} 100{100} - 5.7{5.7} $3656\text{ (Model, Grad., Prompt)}$}
FedOne-BBT 5036.15{5036.15} 0.38{0.38} 100{100} 20000{20000} $1.9\times 10^{-3}$} $1.90\times 10^{-3}\text{ (Prompt Only)}$}
FedOne-BDPL 234.92{234.92} 3.05{3.05} 100{100} 2000{2000} $1.5\times 10^{-2}$} $1.52\times 10^{-2}\text{ (Prompt Only)}$}
FedOne-GS-BDPL 234.73{234.73} 3.05{3.05} 100{100} 2000{2000} $1.5\times 10^{-2}$} $1.52\times 10^{-2}\text{ (Prompt Only)}$}
@@
No Prompt ${14.05}_{0.00}$ ${68.04}_{0.00}$ ${91.35}_{0.00}$ ${79.62}_{0.00}$ ${36.01}_{0.00}$ ${56.22}_{0.00}$ ${72.20}_{0.00}$ 59.64{59.64}
Prompt w/o. Training ${9.17}_{1.31}$ ${62.94}_{4.94}$ ${79.47}_{2.71}$ ${79.88}_{1.94}$ ${25.53}_{1.10}$ ${56.69}_{4.76}$ ${72.16}_{2.51}$ 54.41{54.41}
FedOne-BDPL $13.67_{0.17}$ $68.25_{3.05}$ $87.77_{3.82}$ $81.69_{0.70}$ $32.16_{6.51}$ $0.58_{1.35} 77$$7.62_{1.62}$ 61.67
@@ A.1 Notations and Omitted Mathematical Steps
t=0,,T1t=0,...,T-1 Number of global iterations
e=1,,Ee=1,...,E Number of local iterations
s=0,,Ss=0,...,S Number of client-server interactions
$\left\|\cdot\right\|$ The default 2-norm in this paper
@@
$\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\cdot,\cdot)$ The local stochastic mini-batch variance-reduced policy gradient of the client $k$ (2.2)
@@
$\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{\ast}(\cdot,\cdot)$ The average stochastic mini-batch variance-reduced policy gradient of the client group $U_{t}$ (92)
@@
For i=1,,ni=1,...,n, we define the stochastic gradient, stochastic mini-batch gradient and full gradient with respect to $\boldsymbol{\alpha}^{k}_{i}$ as following:
@@ A.3 Lemmas
Unbiasedness and bounded variance of variance-reduced policy gradient: At the $k$-th client, $\forall\ \psi_{m}^{k}\in\Psi^{k}$, $r=1,\cdots,I$ denotes the $r$-th sampling of $\Phi^{k}$ from $\boldsymbol{p}^{k}$ w.r.t. $\left\{\Phi^{k,r}\right\}_{r=1}^{I}\sim\boldsymbol{p}^{k}$, $\boldsymbol{p}^{k}=\text{GS}(\boldsymbol{\alpha}^{k})$, $\alpha_{i,j}\geq\nu>0$ for i=1,,ni=1,...,n and j=1,,Nj=1,...,N, $\tau>0$ is the temperature parameter, and $\sigma_{\alpha}^{2}=\frac{8G^{2}N}{\tau^{2}\nu^{2}}$, then the variance-reduced policy gradient is unbiased and bounded by :
@@
$\overset{(1)}{=}\frac{1}{I}\sum_{r=1}^{I}\mathbb{E}_{\Phi^{k,r}}\left[\mathcal{L}(\Phi^{k,r},\psi_{m}^{k})\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}\log P(\phi_{i}^{k,r})\right]$
@@
$\overset{(2)}{=}\nabla_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k},\psi_{m}^{k})-\frac{1}{I}\sum_{r=1}^{I}\left(\frac{1}{I-1}\sum_{\begin{array}[]{c}w=1\\ w\neq r\end{array}}^{I}\mathbb{E}_{\Phi^{w}}\left[{\mathcal{L}}(\Phi^{k,w},\psi_{m}^{k})\right]\right)\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}\left(\sum_{\phi_{i}^{k,r}\sim\boldsymbol{p}^{k}_{i}}P(\phi_{i}^{k,r})\right)$ (46)
$\overset{(3)}{=}\nabla_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k},\psi_{m}^{k})-\frac{1}{I}\sum_{r=1}^{I}\left(\frac{1}{I-1}\sum_{\begin{array}[]{c}w=1\\ w\neq r\end{array}}^{I}\mathbb{E}_{\Phi^{w}}\left[{\mathcal{L}}(\Phi^{k,w},\psi_{m}^{k})\right]\right)\cdot\nabla_{\boldsymbol{\alpha}^{k}_{i}}\left(1\right)$ (49)
@@
$\overset{(1)}{=}\frac{1}{I^{2}}\sum_{r=1}^{I}\mathbb{E}_{\Phi^{k,r}}\left[\left\|\frac{1}{I-1}\sum_{\begin{array}[]{c}w=1\\ w\neq r\end{array}}^{I}\left(\mathcal{L}(\Phi^{k,r},\psi_{m}^{k})-{\mathcal{L}}(\Phi^{k,w},\psi_{m}^{k})\right)\nabla_{\boldsymbol{\alpha}^{k}_{i}}\log P(\phi_{i}^{k,r})-\nabla_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k},\psi_{m}^{k})\right\|^{2}\right]$ (55)
@@
L-smooth for $\boldsymbol{\alpha}^{k}$: At the $k$-th client, the $\Phi^{k}$ is sampled from probability matrix $\boldsymbol{p}^{k}$, and $\boldsymbol{p}^{k}=\text{GS}(\boldsymbol{\alpha}^{k})$, $\alpha_{i,j}\geq\nu>0$ for i=1,,ni=1,...,n and j=1,,Nj=1,...,N, $\tau>0$ is the temperature parameter, $\mathbb{E}_{\Phi^{k}\sim\text{GS}(\boldsymbol{\alpha}^{k})}\left[{\mathcal{L}}(\Phi^{k},\Psi^{k})\right]$ is L-smooth for $\boldsymbol{\alpha}^{k}$ and $L=\frac{nGN(\tau+1)}{\tau^{2}\nu^{2}}$, and then for $t$-th iteration, the following inequality is satisfied:
@@
$\leq\left<\nabla_{\boldsymbol{\alpha}^{k}}\mathbb{E}_{\Phi^{k}_{t}\sim\text{GS}(\boldsymbol{\alpha}^{k}_{t})}\left[{\mathcal{L}}(\Phi^{k}_{t},\Psi^{k})\right],\boldsymbol{\alpha}_{t+1}^{k}-\boldsymbol{\alpha}_{t}^{k}\right>+\frac{L}{2}\left\|\boldsymbol{\alpha}_{t+1}^{k}-\boldsymbol{\alpha}_{t}^{k}\right\|^{2}$
@@
[…]
@@
Convergence of BDPL: Suppose Assumption 3.1 and 3.2 hold, for t=0,,T1t=0,...,T-1, $\alpha_{i,j}\geq\nu>0$ for i=1,,ni=1,...,n and j=1,,Nj=1,...,N, $\tau>0$ is the temperature parameter, $\mathbb{E}_{\Phi^{k}\sim\text{GS}(\boldsymbol{\alpha}^{k})}\left[{\mathcal{L}}(\Phi^{k},\Psi^{k})\right]$ is L-smooth for $\boldsymbol{\alpha}^{k}$ and $L=\frac{nGN(\tau+1)}{\tau^{2}\nu^{2}}$, $\sigma_{\psi}^{2}$ is the variance of the stochastic gradient, $\sigma_{\alpha}^{2}=\frac{8G^{2}N}{\tau^{2}\nu^{2}}$ is the variance of the variance-reduced policy gradient. Let $\eta\leq\frac{1}{L}$, then the BDPL’s full gradient $\nabla F^{k}(\boldsymbol{\alpha}_{t}^{k},\Psi^{k})$ satisfies the following inequality:
@@
$\leq\left<\nabla_{\boldsymbol{\alpha}^{k}}\mathbb{E}_{\Phi_{t}^{k}\sim\text{GS}(\boldsymbol{\alpha}_{t}^{k})}\left[{\mathcal{L}}(\Phi_{t}^{k},\Psi^{k})\right],\boldsymbol{\alpha}_{t+1}^{k}-\boldsymbol{\alpha}_{t}^{k}\right>+\frac{L}{2}\left\|\boldsymbol{\alpha}_{t+1}^{k}-\boldsymbol{\alpha}_{t}^{k}\right\|^{2}$
@@
$=\left<\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{t},\Psi^{k}),-\eta\cdot\mathbb{E}_{\{\Phi^{r}\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}}\left[\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{t},\mathcal{B}^{k}_{t})\right]\right>+\frac{L\eta^{2}}{2}\mathbb{E}_{\{\Phi^{r}\}_{r=1}^{I}}\mathbb{E}_{\mathcal{B}_{t}}\left[\left\|\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{t},\mathcal{B}^{k}_{t})\right\|^{2}\right]$
$\overset{(1)}{=}\left<\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{t},\Psi^{k}),-\eta\cdot\mathbb{E}_{\mathcal{B}_{t}}\mathbb{E}_{\{\Phi^{r}\}_{r=1}^{I}}\left[\hat{\nabla}_{\boldsymbol{\alpha}^{k}_{i}}f^{k}(\boldsymbol{\alpha}^{k}_{t},\mathcal{B}^{k}_{t})\right]\right$>
@@
$\overset{(2)}{=}-\eta\cdot\left<\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{t},\Psi^{k}),\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{t},\Psi^{k})\right>+\frac{L\eta^{2}}{2}\left\|\nabla_{\boldsymbol{\alpha}^{k}_{i}}F^{k}(\boldsymbol{\alpha}^{k}_{t},\Psi^{k})\right\|^{2}$
@@
$\mathbb{E}_{\boldsymbol{u}}\left\|\nabla_{\boldsymbol{\alpha}^{k}}F^{k}(\boldsymbol{\alpha}_{t}^{k},\Psi^{k})\right\|^{2}=\int_{\boldsymbol{0}}^{\boldsymbol{1}}\left\|\nabla_{\boldsymbol{\alpha}^{k}}F^{k}(\boldsymbol{\alpha}_{t}^{k},\Psi^{k})\right\|^{2}\textbf{1}_{n}d\boldsymbol{u}\leq\left\|\nabla_{\boldsymbol{\alpha}^{k}}F^{k}(\boldsymbol{\alpha}_{t}^{k},\Psi^{k})\right\|^{2}$